OpenStudy (mallorysipp234):

this makes no sense. help please? If cos Θ = -2/5 and tan Θ > 0, what is the value of sin Θ?

OpenStudy (anonymous):

\[\sin (\theta )=\sin \left(\cos ^{-1}\left(-\frac{2}{5}\right)\right) \]

OpenStudy (anonymous):

\[\sin (\theta )=\frac{\sqrt{21}}{5} \]

OpenStudy (mallorysipp234):

i got a syntax error on my calculator. im still confused @robtobey

OpenStudy (anonymous):

if cosine is negative, which quadrant do you think theta is in?

OpenStudy (mallorysipp234):

negative square root 21 negative square root 21 / 5 square root 21 / 5 square root 21 / 2 @sourwing @robtobey

OpenStudy (anonymous):

yes, which quadrant theta is in if cosine is negative?

OpenStudy (anonymous):

sin x = -(V21)/5 since tan x > 0. See drawing.

OpenStudy (mallorysipp234):

It isn't square root 21/5

OpenStudy (mallorysipp234):

@thu1935 so it's negative square root 21 / 5?

OpenStudy (anonymous):